jigglypuff314 (jigglypuff314):

Definite Integral problem... Calculate the Average value of the given function of the interval provided. \[f(x) = \frac{ 1 }{ (3 + \sqrt{x})^{4} }~~,~~[4, 9]\]

OpenStudy (shamil98):

\[av(f) = \frac{ 1 }{ b-a } \int\limits_{a}^{b} f(x) dx\]

OpenStudy (anonymous):

do you know the formula?

OpenStudy (anonymous):

hehe ^ like mentioned above :D

OpenStudy (shamil98):

\[\frac{ 1 }{ (3+\sqrt{x})^4 } = (3+\sqrt{x})^{-4}\] use the anti-power rule and the formula ofc.

jigglypuff314 (jigglypuff314):

idk anti power rule, but I do understand the equation stuff you posted... erm... I did u = 3 + rootx \[\int\limits_{5}^{6} 2u ^{-3}-6u ^{-4} du\]???

OpenStudy (shamil98):

you cant use u-sub for this..

jigglypuff314 (jigglypuff314):

:0 why not?

OpenStudy (shamil98):

anti-power rule.. is \[\int\limits_{}^{} x^n = \frac{ x^{n+1} }{ n+1 }\] \[av(f) = \frac{ 1 }{ 5 } \int\limits_{4}^{9} (3+\sqrt{x})^{-4}\] \[= \frac{ 1 }{ 5 } \left[ -\frac{ 1 }{ 3(3+\sqrt{x})^3 } \right]^9_4\]

OpenStudy (shamil98):

just evaluate that using the fundamental theorem of calc now. and simplify.

jigglypuff314 (jigglypuff314):

nuh uh me didn't learn dat

OpenStudy (shamil98):

and you're doing u-sub??

jigglypuff314 (jigglypuff314):

yeah...

OpenStudy (shamil98):

techniques of integration are AFTER the basic rules -.-

jigglypuff314 (jigglypuff314):

meh teacher dumbdumb

jigglypuff314 (jigglypuff314):

or teacher went over anti power rule, but not tell us name... could you give simple example of anti power rule?

jigglypuff314 (jigglypuff314):

wut happened to chain rule in the thing you gave?

OpenStudy (shamil98):

\[\int\limits_{4}^{9} \frac{ 1 }{ (3+\sqrt{x})^4 } dx\] even if you let \[\ u = 3 + \sqrt{x} \] your \[\ du = \frac {1}{2\sqrt{x}} dx \] \[\ 2~ du = \frac {dx}{\sqrt{x}}\] you can't do u-sub cause you don't have the 1/(sqrt{x}) needed...

jigglypuff314 (jigglypuff314):

put du so you have 2rootx to even it out then rootx = u-3

OpenStudy (shamil98):

i'd rather you try to learn this stuff from the beginning starting with the basic rules of integration.. i don't know why you're doing attempting u-sub without prior knowledge.

OpenStudy (shamil98):

and that won't work.

OpenStudy (shamil98):

http://www.mathsisfun.com/calculus/integration-rules.html there's a list of the rules.

OpenStudy (anonymous):

you can try u = sqrt(x) x = u^2 dx = 2u du